OpenStudy (anonymous):

y=ab^x (-3,24)(-2,12) exponetial equation

OpenStudy (watchmath):

\(24=a\cdot b^{-3}\) \(12=a\cdot b^{-2}\) Eq 2 /eq 1 1/2 = b Plug in to the second equation \(12=a/4\) \(a=48\)

OpenStudy (watchmath):

Sorry 1/2 = a/4 so a = 2

OpenStudy (watchmath):

sorry again... the first one is correc :D :D :D

OpenStudy (anonymous):

the answer choices r -3(24)^x b. y=3x y=3(0.5)^x y=0.5(3)^x

OpenStudy (watchmath):

4a = 12 so a=3

OpenStudy (anonymous):

sry should of specified it

OpenStudy (watchmath):

so y=(0.5)^x

OpenStudy (watchmath):

y=3(0.5)^x

OpenStudy (anonymous):

thx do u want 2 help with more questions lol?

OpenStudy (watchmath):

Just post your questions. There are lots of people that willing to help :D.